OpenStudy (anonymous):

What is the axis of symmetry of the graph of y = 2(x – 3)2 – 7? A)x = –6 B) x = 2 C)x = 3 D)x = 7

OpenStudy (anonymous):

Find the derivative. set it to zero and solve for x.

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

multiply out y = 2x^2 -12x + 11

OpenStudy (anonymous):

vertex is at -b/2a

OpenStudy (anonymous):

so the answer would be A

OpenStudy (anonymous):

the sign is wrong

OpenStudy (anonymous):

-b/2a = -(-12/4) = 3

OpenStudy (anonymous):

you lost me im confused again

OpenStudy (anonymous):

we obtained an equation y = 2x^2 -12x + 11 there is a formula to find the vertex x = -b/2a do you know it?

OpenStudy (callisto):

For a quadratic equation in the vertex form, y = a(x-h)^2 + k axis of symmetry is x = h while the vertex is at (h,k) Now, you have f y = 2(x – 3)^2 – 7, identify what h is in this case and you can get your answer.

OpenStudy (anonymous):

so you have the same answer!

OpenStudy (anonymous):

when i do that i get 3 so that would be the answer correct?